One last darn problem..
A square: the top is labeled b+4, the right side is labeled, b.
The perimeter is?
The area is?
:-)

Answer by rfer(16322)

P=2(b+4)+2b
P=2b+8+2b
P=4b+8
-----------------
A=(b+4)(b)
A=b^2+4b